Imagine the market for the phone card has a demand function of QDX=38-2px and the supply function of QSX=4px-10wheres px is the price of the phone card.if px=10 calculate the producer surplus PS.


1
Expert's answer
2021-03-08T07:18:19-0500

Solution:

Find the equilibrium: QDX = QSX

38 – 2px = 4px – 10

38 + 10 = 4px + 2px

48 = 6px


px = "\frac{48}{6}" = 8


Equilibrium price (px) = 8

Derive quantity:

38 – 2px = 38 – "(2\times 8)" = 38 – 16 = 22

Equilibrium quantity = 22

Market price given = 10

Producer surplus ="(Market \;price - Minimum\;price)\times Quantity\; sold"

="(10 - 8)\times 22"

= "2\times22 = 44"


Producer Surplus = 44